Comment “Fatty acids are storage fuels”. The weight of a person ABC is 65 kg. About 11.5 kg of his weight is due to the fat deposits in the body. He went for a hunger strike. Can you explain how long he can survive? i. without water and food? ii. without food but with water? 


1
Expert's answer
2021-03-30T04:55:42-0400

P = 370 + (21.6 * (65 kg - 11.5 kg)) = 1525.6 kcal/day this person need

1g of fat = 9kcal

11.5 kg = 11500g

11500 g * 9 kcal /g= 103500 kcal

103500 kcal / 1525.6 kcal/day = 67.8 day person can live without food

For the hydrolysis of fat, a person needs water, so a person cannot live longer than a week without water.